John sells cookies for $1.5 apiece. Each day he sells 77 cookies. The cost of making each cookie is $0.5. What is John's daily revenue? (Round to the nearest whole number)
Answer #1

Total Revenue= P*Q = $1.5 * 77 = $115.5 is approximately$116

Total Profit= P*Q - C*Q = $1.5*77 - $0.5

= $77